In a company having 180 employees, 60 earn $10.00 per hour and 120 earn $12.00 per hour. 30 employees each earns $12.00 leave the work. Determine the mean earnings per hour.


1
Expert's answer
2022-03-29T17:12:26-0400

"EX=\frac{10\cdot 60+12\cdot 120}{180}=11.3333"